Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

29. Total number of balls =18

Number of red balls =8

Number of black balls =10

i. Probability of getting red ball at first draw =818=49

The red ball is replaced after the first draw

∴ Probability of getting red ball in second draw =818=49

∴ Probability of getting both the balls red =49×49=1681

ii. Probability o

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

28. The sample space of given condition is

S= {1, 2, 3, 4, 5, 6}

Let,  P (A)= probability of getting an odd number in first throw

⇒P (A)=36=12

P (B)= probability of getting an even number

⇒P (B)=36=12

Probability of getting an even number in three times =12×12×12=18

So, Probability of getting an odd number at least once

=1− probability of getting an odd

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

27. Given,

P(A)=0.3

P(B)=0.6

i. A and B are independent

⇒P(A∩B)=P(A).P(B)

=0.3×0.6

=0.18

ii. P ( A and not B )

⇒P(A∩B′)=P(A)−P(A∩B)

=0.3−0.18

=0.12

iii. P ( A or B )

⇒P(A∪B)=P(A)+P(B)−P(A∩B)

=0.3+0.6−0.18

=0.72

iv. P (neither A nor B )

⇒P(A′∩B′)=P(A∪B)′

=1−P(A∪B)

=1−0.72

=0.28

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

26. Given,

P(A)=12

P(B)=712

P (not A or not B )= 14

⇒P(A′∪B′)=P(A∩B)′

⇒P(A∩B)′=14

⇒1−P(A∩B)=14

⇒P(A∩B)=1−14

⇒4−14

⇒34

Now,

P(A).P(B)=12×712

=724≠34

⇒P(A).P(B)≠P(A∩B)

Therefore, A and B are not independent.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

25. Given,

P (A)=14

P (B)=12

P (A∩B)=18 Here,  P  (not A not B ) =P (A′∩B′)=P (A∪B)′

⇒1−P (A∪B)

⇒1− [P (A)+P (B)−P (A∩B)]

⇒1− [14+12−18]

⇒1− [2+4−18]

⇒1−58=8−58

⇒38

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

24. Given,

P(A)=0.3

P(B)=0.4

A and B are independent

(i) P(A∩B)=P(A).P(B)

=0.3×0.4=0.12

(ii) P(A∪B)=P(A)+P(B)−P(A∩B)

P(A∪B)=0.3+0.4−0.12=0.58

(iii) P(A/B)=P(A∩B)P(B)=0.120.4=0.3

(iv) P(B/A)=P(B∩A)P(A)=0.120.3=0.4

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

23. Given,

P(A)=12

P(A∪B)=35

P(B)=p

(i) We know that,

when A and B are mutually exclusive

(A∩B)=∅ ,P(A∩B)=0

P(A∪B)=P(A)+P(B)−P(A∩B)

⇒35=12+p

⇒p=35−12=6−510=110

(ii) when A and B are independent

P(A∩B)=P(A).P(B)P(A∩B)=12×p

Now,

P(A∪B)=P(A)+P(B)−P(A∩B)

⇒35=12+p−12p

⇒35−12=2p−p2

⇒6−510=p2

⇒p=1×210=210=15

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

22 . P (F)=310

P (E∩F)=15

P (E).P (F)=35×310

=950≠P (E∩F)

Here,  P (E).P (F)≠P (E∩F) . A and B are not independent.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

21. When die is thrown, the sample space

S= {1, 2, 3, 4, 5, 6}

Let,  A =the number is even = {2, 4, 6}

B =the number is red = {1, 2, 3}

P (A)=36=12P (B)=36=12A∩B= {2}P (A∩B)=16∴P (A).P (B)=12×12=14≠P (A∩B)

Hence,  A and B are not independent.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

20. The sample space of given condition are

S={(H,1),(H,2),(H,3),(H,4),(H,5),(H,6)(T,1),(T,2),(T,3),(T,4),(T,5),(T,6)}

Let A = head appear on the coin

B = 3 on the die

A={(H,1),(H,2),(H,3),(H,4),(H,5),(H,6)}

P(A)=612=12

B={(H,3),(T,3)}

P(B)=212=16

A∩B={(H,3)}

i.e.,P(A∩B)=112

P(A).P(B)=12×16=112=P(A∩B)

Therefore, A and B are independent.

Thumbs Up IconUpvote Thumbs Down Icon